# Multi-Class Stackelberg Games for the Co-Design of Networked Systems

### v1

@@ VI Concluding Remarks and Future Directions
We have presented multiple classes of Stackelberg games for the co-design of networked systems comprising the simultaneous design of both system and control. Under this approach, a leader is in charge of deciding on a design system parameter. Note that, in general, this decision can be related to the selection of elements such as actuators, or any other system specification. Then, there is a follower layer where the control design takes place. As shown in this paper, the decisions made at the control design depend on the decisions made at the system design stage. Hence, we have presented multiple possibilities for such a bi-level Stackelberg-like interaction. This is because we can consider multiple parties at each one of the layers, i.e., multiple leaders and multiple followers, leading to more involved game-theoretical settings. We consider the case in which leaders and followers can either cooperate or not, and all the possible combinations for these interactions. We have shown that the cooperative game approach coincides with a control problem, and the non-cooperative game can be seen as a decentralized control strategy. Moreover, the evaluation of the price-of-anarchy for the computed equilibrium solutions shows that the decentralized controllers are optimal. Finally, as it was highlighted in the manuscript, we have considered a discrete finite set of strategies for the leaders. As future work, it is proposed to consider a continuum set for the strategic leader set.
Acknowledgments
Julian Barreiro-Gomez is profoundly grateful to God and Our Lady of Lourdes for the blessings of health and life, without whose intercession this work would not have been possible. Ye Wang acknowledges support from the Australian Research Council through the Discovery Early Career Researcher Award (DE220100609).

### v2

@@ VI Concluding Remarks and Future Directions
We have presented multiple classes of Stackelberg games for the co-design of networked systems comprising the simultaneous design of both system and control. Under this approach, a leader is in charge of deciding on a design system parameter. Note that, in general, this decision can be related to the selection of elements such as actuators, or any other system specification. Then, there is a follower layer where the control design takes place. As shown in this paper, the decisions made at the control design depend on the decisions made at the system design stage. Hence, we have presented multiple possibilities for such a bi-level Stackelberg-like interaction. This is because we can consider multiple parties at each one of the layers, i.e., multiple leaders and multiple followers, leading to more involved game-theoretical settings. We consider the case in which leaders and followers can either cooperate or not, and all the possible combinations for these interactions. We have shown that the cooperative game approach coincides with a control problem, and the non-cooperative game can be seen as a decentralized control strategy. Moreover, the evaluation of the price-of-anarchy for the computed equilibrium solutions shows that the decentralized controllers are optimal. Finally, as it was highlighted in the manuscript, we have considered a discrete finite set of strategies for the leaders.
As future work, we propose extending the framework to consider a continuous strategy space for the leaders. This extension would leverage approximation theory and learning-based methods to address the computational tractability challenges associated with large-scale systems. By adopting these techniques, it becomes feasible to approximate the optimal strategies in high-dimensional settings where exhaustive enumeration is intractable.